OpenStudy (anonymous):

Factor completely: v^4y^4 -16y^4 .

OpenStudy (anonymous):

(v^4-16)y^4

OpenStudy (anonymous):

@jim_thompson5910 is this right factoring isn't usually a problem but I need to make sure lol

OpenStudy (anonymous):

OR is it (V-2)(V+2)(V^2+4)Y^4

jimthompson5910 (jim_thompson5910):

hint: \[\Large v^4 - 16 = (v^2)^2 - (4)^2\]

OpenStudy (anonymous):

What about the y?

jimthompson5910 (jim_thompson5910):

ignore that for now

jimthompson5910 (jim_thompson5910):

difference of squares rule: a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

like in the second equation I did? with the (v-2)(v+2)

jimthompson5910 (jim_thompson5910):

v^4y^4 -16y^4 (v^4 -16)y^4 ((v^2)^2 -16)y^4 ((v^2)^2 -4^2)y^4 (v^2 - 4)(v^2 + 4)y^4 ... apply difference of squares rule (v - 2)(v + 2)(v^2 + 4)y^4 ... apply difference of squares rule again

OpenStudy (anonymous):

(V-2)(V+2)(V^2+4)Y^4 that's what I put backa t the top =o you're saying I need to apply the rule to it one more time

jimthompson5910 (jim_thompson5910):

yes you'll use the difference of squares rule twice

OpenStudy (anonymous):

ok and this time i'll use it on the y^4?

jimthompson5910 (jim_thompson5910):

no you pull out y^4 and temporarily ignore it while you apply the difference of squares rule

jimthompson5910 (jim_thompson5910):

you only use the difference of squares rule when you have two squares and you're subtracting them

OpenStudy (anonymous):

ugh four more left and my brain stopped working 5 questions ago I even ate salmon! isn't that like the best brain food =p

jimthompson5910 (jim_thompson5910):

why not take a break then

OpenStudy (anonymous):

because I'm soooo close!!! I didn't come this far without a break to stop now it's only 4